UT 20120206 AO/PISCES technical night

Observer: DMiller, JHill (Tucson), JCarlos (LBTO)
Telescope Operator: DHuerta (LBTO)
Additional AO Support: SEsposito, EPinna, APuglisi, MXompero
Telescope Support: JUrban(LBTO)
Instrument Support: DOfficer (LBTO)

Summary

  • Get good images on Pisces after JC and Doug removed condensation from the Pisces window in the afternoon
  • Part way through the night condensation build up on Pisces window again. JC and David use dry air to clear. After this, good images on Pisces.
  • Take two sets of dice5 data, 3 and 6 arcsec to view image quality over Pisces chip
  • Determine offset for PCS to place on-axis target near the center of Pisces chip
  • Determine focus offset for ND3+Ks to be 53.045
  • Arcetri Guys close loop on the SX side while we were waiting for BCU39 to warm up from 4C to 6C
  • Preset in ACE mode to two science objects.
    • ULAS1120: failed becasue of CD39 problems and then, I think, getting too bright as sunrise approach
    • Q1436m0051: Failed because AO Reference start too dim (17.6 R mag). Can this be fixed?
      (JMH: not without going to more than 60 arcsec radius - 14.5 mag star at 133 arcsec distance)
  • We asked Doug Officer to fill Pisces this morning and tonight so JC can take sky flat this evening and maybe get on sky for a bit with Simone to better align the ND holder

Details

Close dome test.

After inspecting the entrance window of the Pisces dewar we saw that it was a condensation in the center of the window. After checking that we don't have condensation a set of images in J were taken with the same camera set up than yesterday.

In the images we can see clearly the differences between the two days.

YESTERDAY... -pisces.20120205.0023 filter J exp time 120" dome lights OFF

# IMAGE MEAN MIDPT STDDEV pisces.20120205.0023.fits[700:800,670:770] 42.8 42.62 11.47

-pisces.20120205.0021 filter J exp time 120" dome lights ON

# IMAGE MEAN MIDPT STDDEV pisces.20120205.0021.fits[700:800,670:770] 4463. 4456. 291.4

TODAY AFTER REMOVE THE CONDENSATION.

pisces.20120205.0091 filter J, exptime 120" dome lights OFF

# IMAGE MEAN MIDPT STDDEV pisces.20120205.0091.fits[700:800,670:770] 252.4 252.6 18.14

pisces.20120205.0089 filter J, exptime 120" dome lights ON... WELL SATURATED IMAGE

# IMAGE MEAN MIDPT STDDEV pisces.20120205.0089.fits[700:800,670:770] -26906. -28704. 10410.

Handover

First half of night was TCS testing of Build 2012B, although most of the night was consumed by UPS and server restarting.

7:10 David reconfigure telescope for IRTC@RFBG when we are back using TCS build BP14.

7:20 JC, Simone go up to Pisces to install ND holder. The holder has a target on it and they shone a flashlight on it. Enrico was taking image and instructing moves.
  • Could see an edge on upper right
  • Move and now constant illumination on Pisces
  • images 5-8

  • Procedure of using a pinhole willbe done tomorrow

7:40 David is setting pointing and collimate on BS star

8:00 Preset to AO302 in Active mode. WFS has been powered off and is too cold (the BCU analog circuit that drives the camera lens).

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO302
preset, OBJECT=AO302, MODE=active, GS=2, PA=39
auto_collimate

  • 6.8 R mag
  • 5.0 H mag

8:20 Find on-axis position

  • images 10-12
  • star at 200, 523
  • move to near center on upper left quadrant

8:30 iif_offset, 1, -5, COORD='DETXY'

  • Put star on pixel 460, 570
  • Call this center
  • Absorb x=1, y=-5
  • PCSInstrument x=0.6, y = -3.0

8:40 John find SX M1 and M2 offsets in coma-free pointing to find on-axis star on LBBG which is already acquired on Right side.

  • tip = -45 arcsec
  • tilt = +49 arcsec

JMH: The following night, the converse offsets in coma-free pointing to find the on-axis star on RFBG which is already acquired on Left side.

  • tip = +42 arcsec
  • tilt = -52 arcsec

8:45 observe, dither.obs, /OPEN

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO302
camera, EXP=40.0, FRAMES=1, COADDS=1, FILTER=4
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=5.0, ITER=1, IMAGE=SKY

  • images 19-23
  • lower left and upper right images have much less flux...why

8:55 iif_offset, 5, -5, COORD='DETXY'

  • image 24: bad readout
  • images 25-26 upper right dim star

9:03 iif_offset, 2, -2, COORD='DETXY'

  • image 27: still dim

9:03 iif_offset, 0, 0, COORD='DETXY'

9:05 BCU47 controller is too cold (4 C) needs to be 6 C. If below 6 C then get crosstalk between X and Y and lens is unstable. Waiting...
Arcetri guys are working on left side (authorized on IIF as None).

9:11 AO guys close the AO loop on the left side. 500 modes!!

9:45 test of getimage observe command. image 29

9:49 BCU47 at 5.5 C

9:50 test of getimage observe command, image 30

10:20 BCU47 is now 6 C

observe, 'script./preset.obs'

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO302
preset, OBJECT=AO302, MODE=ace, GS=2, RS=0, PA=40
auto_collimate
runao

  • Failed. A NFS mounted directory was not available. Not properly mounted after shutdown and startup this evening

===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = AcquireRef[13] complete -  dx:-0.26 dy:0.324 slNull: f1:Filter #1 f2:Filter #0 freq:990 mag:7.71 nBins:1 nModes:400 r0:0 snMode:0 strehl:0 ttMod:3
        err = Resmsg[2]  = [Errno 2] No such file or directory: '/towerdata/adsec_calib/M2C/KL_v7/gain/g_tt_0.20_ho1_0.20_ho2_0.20.fits'
        err = Resmsg[3]  = StartAO[14] failed - [Errno 2] No such file or directory: '/towerdata/adsec_calib/M2C/KL_v7/gain/g_tt_0.20_ho1_0.20_ho2_0.20.fits'

10:26 observe, 'script./preset.obs'

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO302
preset, OBJECT=AO302, MODE=ace, GS=2, RS=0, PA=40
auto_collimate
runao

  • Success

10:35 Take images 35-38. Very dim

10:42 Change to Ks filter

  • image 40
  • Also very dim
  • shaped like a galaxy, not a star

10:47 JC run focus routine. images 40-52

  • Ks best focus at 53.051 mm
  • PSF is poor

10:59 offset, 6, 0, COORD='DETXY'

11:01 observe, preset.obs

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO302
preset, OBJECT=AO459, MODE=ace, GS=2, RS=0, PA=152
auto_collimate
runao

  • first guide images look like a 2x2 box
  • AO Closed loop
  • pisces image 54, very dim and looks like a galaxy. CCD47 image looks diffraction limited
  • offset, -6, 0
  • image 55, very dim and looks like a galaxy. CCD47 image looks diffraction limited

11:26 Telescope at Zenith. JC and David go to take a look at the Pisces window.

12:00 JC and David report a lot of condensation on the Pisces window

Image of the condensation that cover most of the entrance window. condensation_20120206.JPG

12:02 observe, preset.obs

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO302
preset, OBJECT=AO459, MODE=ace, GS=2, RS=0, PA=112
auto_collimate
runao

  • image 66: much brighter, saturated
  • shorter exposures, still saturated images 68
  • go to star 2 mag dimmer

12:14

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO442
preset, OBJECT=AO442, MODE=ace, GS=1, RS=0, PA=29
auto_collimate
runao

  • image 69, 70. Near quadrant edge. Move down 2"
  • image 71 Good. 5000 peak counts

Optical loop diagnostic

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120206_121048 ??? 7.5 70 12 NaN 0.53 092604 1 400 990 0.7 0.9 0.7 3 585 ON 65.7 Ks 1800 10 ONSKY

12:28 JC start focus routine

Focus determined for: ND3.0+Ks is 53.045

12:40 Start dither pattern with 3 and 6 arcsec dice5. I didn't set_position. Kill

12:46 observe, dither.obs

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO302
set_position
camera, EXP=1.0, FRAMES=1, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=3.0, ITER=1, IMAGE=SKY
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=6.0, ITER=1, IMAGE=SKY

  • start image 90

Dither
X Y frames Exposure Filter PISECS Image TYPE SR
0.00000 0.00000 1 of 1 1.00000 Ks pisces.20120206.10090.fits SKY 69
3.00000 3.00000 1 of 1 1.00000 Ks pisces.20120206.10091.fits SKY 49
-3.00000 3.00000 1 of 1 1.00000 Ks pisces.20120206.10092.fits SKY 67
-3.00000 -3.00000 1 of 1 1.00000 Ks pisces.20120206.10093.fits SKY 74
3.00000 -3.00000 1 of 1 1.00000 Ks pisces.20120206.10094.fits SKY 68
Dither
X Y frames Exposure Filter PISECS Image TYPE SR
0.00000 0.00000 1 of 1 1.00000 Ks pisces.20120206.10095.fits SKY 60
6.00000 6.00000 1 of 1 1.00000 Ks pisces.20120206.10096.fits SKY 62
-6.00000 6.00000 1 of 1 1.00000 Ks pisces.20120206.10097.fits SKY na
-6.00000 -6.00000 1 of 1 1.00000 Ks pisces.20120206.10098.fits SKY 70
6.00000 -6.00000 1 of 1 1.00000 Ks pisces.20120206.10099.fits SKY 64

13:04 observe, ULAS1120

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=ULAS1120
preset, OBJECT=Feb12_ULAS1120, MODE=ace, GS=3, RS=0, PA=40
auto_collimate
runao

  • Incorrect reference star. Should be 1. presetAO correctly failed because Ref star too bright (0.0)

13:06 observe, ULAS1120

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=ULAS1120
preset, OBJECT=Feb12_ULAS1120, MODE=ace, GS=3, RS=1, PA=40
auto_collimate
runao

13:15 CCD39 not responding. JC reset

13:06 observe, ULAS1120

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=ULAS1120
preset, OBJECT=Feb12_ULAS1120, MODE=ace, GS=3, RS=1, PA=40
auto_collimate
runao

  • See something faint on CCD47. image 100, 101
       runao: closing the AO loop
===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = RETRY: No stars found in TV frame (-20001) WFSARB_STAR_NOTFOUND
        err = Resmsg[2]  = AcquireRef[132] failed - RETRY: No stars found in TV frame (-20001) WFSARB_STAR_NOTFOUND

13:27 observe, Q1436m0051

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=Q1436m0051
preset, OBJECT=Feb12_Q1436m0051, MODE=ace, GS=3, RS=2, PA=-20
auto_collimate
runao
runao
closing the AO loop
===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = RETRY: No stars found in TV frame (-20001) WFSARB_STAR_NOTFOUND
        err = Resmsg[2]  = AcquireRef[132] failed - RETRY: No stars found in TV frame (-20001) WFSARB_STAR_NOTFOUND

Darks

exptime(sec) frames file
1 10 pisces.20120206.113 to 121
3 10 pisces.20120206.132 to 141
10 10 pisces.20120206.122 to 131
60 10 pisces.20120206.142 to 151

-- JuanGuerra - 06 Feb 2012
  • Screenshot of Close Loop Bin1:
    opticalloop_bin1.png

  • Modal plot of loop performances:
    modalplot_20120206_b1_0.png

  • Condensation_PSF_close loop:
    CONDENSATION_psf.gif

  • After cleaning the entrance window:
    after_cond20120206-2.png

  • Ice concentrate in most of the entrance window:
    condensation_20120206.JPG />
Topic attachments
I Attachment Action Size Date Who Comment
CONDENSATION_psf.gifgif CONDENSATION_psf.gif manage 130 K 06 Feb 2012 - 23:45 UnknownUser Condensation_PSF_close loop
after_cond20120206-2.pngpng after_cond20120206-2.png manage 131 K 08 Feb 2012 - 12:51 UnknownUser After cleaning the entrance window
condensation_20120206.JPGJPG condensation_20120206.JPG manage 1 MB 10 Feb 2012 - 18:08 UnknownUser Ice concentrate in most of the entrance window
modalplot_20120206_b1_0.pngpng modalplot_20120206_b1_0.png manage 14 K 06 Feb 2012 - 14:04 UnknownUser Modal plot of loop performances
opticalloop_bin1.pngpng opticalloop_bin1.png manage 523 K 06 Feb 2012 - 13:56 UnknownUser Screenshot of Close Loop Bin1
Topic revision: r16 - 25 Jul 2014, DouglasFisher
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback